# Solving Copyright Infringement on Short Video Platforms: Novel Datasets and an Audio Restoration Deep Learning Pipeline

### v1

@@ 1 Introduction
Despite their popularity, user-generated content introduces significant copyright challenges. As presented in Figure 1, users frequently overlay arbitrary background music (BGM) on the generic short videos, disrupting the original audio structure and complicating copyright compliance Charfeddine et al. (2022); Chang (2023); Hasanovic et al. (2022); YouTube Help (2023). This practice masks critical audio elements, such as dialogue or original soundtracks (OST), and undermines the accuracy of audio-based near-duplicate video retrieval systems Avgoustinakis et al. (2021), which are essential for identifying infringements. Addressing these issues demands advanced methods capable of removing arbitrary BGM effectively while preserving the integrity of the original audio.
@@
To address this limitation, we integrate Cross-Modal Video-Music Retrieval (CMVMR) technologies Jin et al. (2021); Prétet et al. (2021); Won et al. (2021); Yi et al. (2021) into our pipeline to align restored OSTs with their corresponding video content. Specifically, we fine-tune the UT-CMVMR model Mao et al. (2024), which employs a Video Conformer and Music Conformer enhanced with Optical Flow and Rhythm Quantization to extract spatial-temporal and rhythmic features. Using cross-attention mechanisms and optimizing with Triplet Loss, the model enhances the similarity between video and music embeddings without relying on metadata or tags, making it ideal for short-video environments.
@@
This study introduces a novel and optimized pipeline for short video platforms like YouTube Shorts by integrating advanced audio separation and cross-modal retrieval techniques. Our approach effectively removes arbitrary BGM and restores OST integrity, addressing critical copyright challenges. The pipeline is rigorously trained and evaluated using our domain-specific datasets, OASD-20K and OSVAR-160, setting a new benchmark for arbitrary BGM removal in short video restoration tasks. The primary contributions are as follows:
@@ 4.0.1 Initial Audio Separation: Background Music and Speech Extraction
The first stage of the pipeline focuses on separating mixed audio (arbitrary BGM and original OST) and speech components from the video’s audio track, providing clean inputs for downstream tasks. This stage consists of two key modules: (1) Background Music Separator: The Background Music Separator utilizes the U-Net–based Demucs model Défossez et al. (2019) to decompose audio into its constituent tracks, such as vocals, bass, and drums. Demucs operates in the time domain, leveraging recurrent layers to capture long-term dependencies Ronneberger et al. (2015), which ensures high-quality separation. The separated BGM components are then consolidated into a single “Mixed-Music” track using Pydub, creating a unified input for subsequent modules.
@@ 6.0.2 Video-Music Matching
UT-CMVMR’s strong performance can be attributed to its architectural innovations, such as the integration of Optical Flow and Rhythm Quantization modules, which enhance its ability to capture temporal dynamics and rhythmic patterns. The use of advanced cross-modal attention mechanisms further facilitates effective interaction between video and audio embeddings, enabling precise alignment. Fine-tuning on the domain-specific HIMV-200K dataset equips the model to address the unique challenges of short video platforms, ensuring it can recommend rhythmically and contextually appropriate audio.
@@ 6.0.3 Pipeline Performance
When comparing arbitrary BGM separation to the dialogue-only audio scenario (which achieves an average SDR of 131.24 dB and SI-SDR of 14.00 dB), ours demonstrates slightly better performance. This showcases the model’s ability not only to effectively remove the BGM but also to reconstruct the OST in a manner that closely matches the original audio, preserving essential spectral and temporal details.

### v2

@@ 1 Introduction
Despite their popularity, user-generated content introduces significant copyright challenges. As presented in Figure 1, users frequently overlay arbitrary background music (BGM) on the generic short videos, disrupting the original audio structure and complicating copyright compliance Charfeddine et al. (2022); Hasanovic et al. (2022); YouTube Help (2023). This practice masks critical audio elements, such as dialogue or original soundtracks (OST), and undermines the accuracy of audio-based near-duplicate video retrieval systems Avgoustinakis et al. (2021), which are essential for identifying infringements. Addressing these issues demands advanced methods capable of removing arbitrary BGM effectively while preserving the integrity of the original audio.
@@
To address this limitation, we integrate Cross-Modal Video-Music Retrieval (CMVMR) technologies Jin et al. (2021); Prétet et al. (2021); Won et al. (2021); Yi et al. (2021) into our pipeline to align restored OSTs with their corresponding video content. Specifically, we fine-tune the UT-CMVMR model Mao et al. (2024), which uses a Video Conformer and Music Conformer enhanced with Optical Flow and Rhythm Quantization to extract spatial-temporal and rhythmic features. Using cross-attention mechanisms and optimizing with Triplet Loss, the model enhances the similarity between video and music embeddings without relying on metadata or tags, making it ideal for short-video environments.
@@
This study introduces a novel and optimized pipeline for short video platforms like YouTube Shorts by integrating advanced audio separation and cross-modal retrieval techniques. Our approach effectively removes arbitrary BGM and restores OST integrity, addressing critical copyright challenges. The pipeline is rigorously trained and evaluated using our domain-specific datasets, OASD-20K and OSVAR-160, setting a new benchmark for arbitrary BGM removal in short video restoration tasks. The contributions are as follows:
@@ 4.0.1 Initial Audio Separation: Background Music and Speech Extraction
The first stage of the pipeline focuses on separating mixed audio (arbitrary BGM and original OST) and speech components from the video’s audio track, providing clean inputs for downstream tasks. This stage consists of two key modules:
(1) Background Music Separator: The Background Music Separator utilizes the U-Net–based Demucs model Défossez et al. (2019) to decompose audio into its constituent tracks, such as vocals, bass, and drums. Demucs operates in the time domain, leveraging recurrent layers to capture long-term dependencies Ronneberger et al. (2015), which ensures high-quality separation. The separated BGM components are then consolidated into a single “Mixed-Music” track using Pydub, creating a unified input for subsequent modules.
@@ 6.0.2 Video-Music Matching
UT-CMVMR’s strong performance can be attributed to its architectural innovations, such as the integration of Optical Flow and Rhythm Quantization modules, which enhance its ability to capture temporal dynamics and rhythmic patterns. The use of advanced cross-modal attention mechanisms further facilitates effective interaction between video and audio embeddings, enabling precise alignment. Fine-tuning on the domain-specific HIMV-200K dataset equips the model to effectively address the unique challenges of short video platforms, ensuring it can recommend rhythmically and contextually appropriate audio.
@@ 6.0.3 Pipeline Performance
When comparing arbitrary BGM separation to the dialogue-only audio scenario (which achieves an average SDR of 131.24 dB and SI-SDR of 14.00 dB), ours demonstrates slightly better performance. This showcases the model’s ability not only to effectively remove the BGM but also to reconstruct the OST in a manner that closely matches the original audio, significantly preserving essential spectral and temporal details.
@@ 7 Conclusion
Acknowledgements
This research was supported by the MSIT (Ministry of Science and ICT), Korea, under the ICAN (ICT Challenge and Advanced Network of HRD) support program (RS-2024-00436934), and the ITRC (Information Technology Research Center) program (IITP-2025-RS-2024-00436936) supervised by the IITP (Institute for Information & Communications Technology Planning & Evaluation). Moreover, this work was supported by the IITP grant funded by the MSIT (RS-2024-00439139, Development of a Cyber Crisis Response and Resilience Test Evaluation Systems).